At what annual rate would the following have to be invested?
$9,337, to grow to $55,466, in 5 years.

Round the answer to two decimal places in percentage form
Answer #1

Ans : Formula to compute future value of sum invested at a rate of r for n number of years compounded yearly ,

FV = PV * (1+r)n

FV = Future value = $ 55,466

PV = Present value = $ 9,337

r = rate of interest = ?

n = number of years = 5 years

Insert above data in the formula ,

55,466 = 9337 * (1+r)5

(1+r)5 = 55466 / 9337

(1+r)5 = 5.9404519

(1+r) = (5.9404519)(1/5)

(1+r) = 1.428117

r = 42.81 %